OpenStudy (anonymous):

the total score in a football game was 72 points. the winning team scored 12 more points than the losing team. how many points did the winning team score?

OpenStudy (ddcamp):

If the winning team got A points and the losing team got B points, we know that: A + B = 72 (The total score) A = B + 12 (The winning team got 12 more points) Can you solve that system of equations?